Please help me solve this problem... i need it rn
When john and marsha worked together, they finished a job in 10 days. But when she joined him only after he had worked for 7 days, they finished the same job in 5 more days. In how many days can each of them finish the job alone?
a.Representation

B.formulate the equation
C. Show the solution
D. Checking

Answer by stanbon(75887)

When john and marsha worked together, they finished a job in 10 days. But when she joined him only after he had worked for 7 days, they finished the same job in 5 more days. In how many days can each of them finish the job alone?
a.Representation
Together:: time = 10days/job ; rate = 1/10 job/day
John:: time = j days/job ; rate = 1/j job/day
Marsha:: time = m days/job ; rate = 1/m days/job
-------------------
B.formulate the equation
rate + rate = together rate
1/j + 1/m = 1/10
7(1/j) + 5(1/10) = 1
---
7/j = 1/2
j = 14 days/job (John's time to finish alone)
---
(1/14) + 1/m = 1/10
1/m = (1/10) - (1/14)
1/m = (14-10)/140
m = 140/4 = 35 (Marsha's time to finish alone)
-----------
Cheers
stan H.
-----------
C. Show the solution
D. Checking